OpenStudy (anonymous):

The sum of four consecutive integers is -22

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x-1)+x+(x+1)+(x+2)=-22 }\) Where the smallest integer is \(\large\color{slate}{ (x-1) }\)

OpenStudy (anonymous):

Ok so then the four integers would be -1,1,0,1 ?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x-1)+x+(x+1)+(x+2)=-22 }\) \(\large\color{slate}{ x-1+x+x+1+x+2=-22 }\) \(\large\color{slate}{ 4x+2=-22 }\) continue from here

OpenStudy (anonymous):

x=5

OpenStudy (solomonzelman):

you have -22 on the left side, x is not 5, please retry.

OpenStudy (solomonzelman):

I mean -22 on the right (not left) side

OpenStudy (anonymous):

x=-5